# Are You Getting What You Pay For? Auditing Model Substitution in LLM APIs

### v1

@@ 4.1 Text-output-based verification
Table 2: Binary classification accuracy between outputs of quantized and original models on the Ultrachat dataset. Accuracies near 50% indicate failure to differentiate.
Analysis.
Table 2 shows that classification accuracies across all model pairs and embedding models consistently hover around 50. This indicates that regardless of original model series, parameter size, and quantization methods, there are no distinguishing features that allow current state-of-the-art embedding models to perform better than random guessing. This differs significantly from the baseline results reported in Sun et al. (2025), where the same classifier setup achieved approximately 98% accuracy for cross-model classification and around 80% accuracy when comparing larger and smaller models within the same series. Our experiments with different training and testing data configurations yielded similar results, providing strong evidence that text classification approaches are ineffective for identifying quantized model substitution.
4.1.2 Identity prompting (Huang et al., 2025)
Method.
A naive method is to prompt the model to reveal its identity, leveraging the fact that many LLMs are trained to disclose their name or developer. Repeated querying using identity-eliciting prompts (e.g., ”Who are you?”, ”Who developed you?”) followed by simple string matching against known model identifiers has proven effective in classifying models based on their responses (Huang et al., 2025).
Attack: Quantization substitution.
We tested if quantization affects the models’ ability to self-identify accurately. We prompted base and FP8 versions of several models (Llama3.1-70B-Instruct, Gemma2-9b-it, Mistral-7b-v3-Instruct, and Qwen2-72B-Instruct) 500 times each with identity queries at temperature 0.6. We then calculated detection accuracy by string matching for model family names (e.g., ”Llama”), developer information (e.g., ”Meta”).
Attack: System prompt override.
System prompt overrides can circumvent identity-based detection methods. By prepending instructions like “You are model XX,” adversaries can manipulate models to suppress their trained identity information. We evaluated this attack by querying Gemma2-9b-it, Mistral-7b-v3-Instruct, and Qwen2-72B-Instruct about their identity using the same methodology as our previous experiments, but with the system prompt “You are LLaMA, an AI assistant developed by Meta AI.” This approach significantly reduced the accuracy of naive identity prompting across all tested models.
Analysis.
Our results demonstrate consistent limitations in model self-identification. Across all tested models, responses contained only high-level identity information (e.g., family name “Gemma” and developer “Google DeepMind”), lacking specifics like parameter count or quantization status. Direct queries about these technical specifications yielded unreliable results, with success rates below 5% and frequent refusals or hallucinated values. While system prompt overrides generally succeeded, certain models occasionally produced inconsistent responses that contradicted the injected identity (e.g., “I am an AI model trained by Mistral AI, not Meta AI”). These contradictions, along with potential side effects on model behavior for unrelated queries, make quantization-based substitution a more reliable attack. Overall, standard identity prompting proves inadequate for detecting either quantization status or parameter size variations within the same model family.
4.1.3 Model equality testing (Gao et al., 2024)
Method.
Given the assumption that all decoding parameters are the same and we have access to model’s text output only, we follow Gao et al. (2024) to use the Maximum Mean Discrepancy (MMD) to compare two distributions over text completions, $P$ and $Q$. Let $k(\cdot,\cdot)$ be a positive definite kernel function. Then the squared MMD is defined by:
[…]
To tailor the MMD to text data generated by LLM, we will empirically use a Hamming Kernel: $k_{\mathrm{hamming}}(x,x^{\prime})\;=\;\sum_{i=1}^{L}\mathbf{1}\{x_{i}=x^{\prime}_{i}\}$, where $x_{i}$ and $x^{\prime}_{i}$ are the i-th tokens of sample $x$ and $x^{\prime}$, respectively, and $L$ is the the maximum sequence length.
To determine statistical significance after calculating the MMD statistic, we compare the observed MMD value against its null distribution (MMD when $P=Q$). We estimate this null distribution through permutation testing. Specifically, we generate multiple random permutations by repeatedly shuffling samples between the $P$ and $Q$ sets and recalculating the MMD statistic for each permutation. The p-value is then the proportion of permutation MMD values that are greater than or equal to our observed MMD value. We reject the null hypothesis (that $P=Q$) if this p-value is below our chosen significance level $\alpha$. This approach allows us to detect distribution differences with relatively small sample sizes, making it practical for API auditing.
Figure 1: Statistical power of model equality testing for the smaller model under quantization attack and the larger model under substitution attack.
Attack: Randomized model substitution.
For the experiment, we used the completions from a dataset of outputs generated by Llama-3-70B-Instruct and Llama-3.1-70B-Instruct from the INT8 quantized model and the original full-precision model. Throughout all experiments, we maintained identical decoding parameters by using maximum completion length of 50 tokens and temperature $T=1$. The prompt distribution is uniform over 25 random strings sampled from multilingual Wikipedia. For each model, we created mixed distributions with parameter $p\in\{0.0,0.2,0.4,0.6,0.8,1.0\}$, where $p$ represents the proportion of INT8 samples in the mixture. We then evaluated the statistical power of our test, based on the Hamming MMD, to distinguish the mixed output distribution under different parameter $p$ from the original model’s distribution for each of the three models with significance level $\alpha=0.05$. For each comparison, we performed 100 Monte Carlo simulations and used 1000 permutations to estimate p-values, calculating the statistical power as the proportion of simulations that correctly rejected the null hypothesis.
Analysis.
As shown in Figure 1, model equality testing on smaller-sized models has statistical power consistently remains below 0.2 and usually close to 0.05, even without randomized substitution, indicating poor detection capability. Consequently, our random model substitution analysis focuses on larger models where these tests initially demonstrate greater detection capability. When the proportion of substituted samples $p$ reaches 0.2, the test’s statistical power decreases to approximately the significance level $\alpha=0.05$. This indicates that with just 20% substitution, the test’s performance becomes statistically indistinguishable from random chance, showing that random substitution attack is successful in such auditing scenario.
Empirically, the MMD distance between instances of the same model can be quite large even when auditing from reputable API provider like Together.ai [^1] Discussed in Gao et al. (2024). This demonstrates that different implementations of inference frameworks can lead to false test results, indicating that such testing cannot be fully trusted.
4.1.4 Benchmark-based detection (Chen et al., 2023)
Method.
An auditor may deploy benchmark-based detection by querying the API using a series of standardized benchmark tasks with known results. By comparing the model’s predictions against these established baselines on the model of suspect, deviations can indicate that the model is substituted.
@@
Meta-Llama-3-8B-Instruct 62.76 $\pm$ 0.18 65.43 $\pm$ 4.70 20.43 $\pm$ 1.85 22.60 $\pm$ 0.25
Meta-Llama-3-8B-Instruct-FP8 62.30 $\pm$ 0.19 61.64 $\pm$ 0.00 15.24 $\pm$ 0.00 20.20 $\pm$ 0.00
Meta-Llama-3-70B-Instruct 78.17 $\pm$ 0.14 88.36 $\pm$ 1.59 35.57 $\pm$ 1.82 29.55 $\pm$ 0.51
Meta-Llama-3-70B-Instruct-FP8 77.84 $\pm$ 0.15 87.49 $\pm$ 1.97 35.61 $\pm$ 1.83 33.21 $\pm$ 0.25
Gemma-2-9b-it 71.86 $\pm$ 0.12 81.84 $\pm$ 2.13 33.34 $\pm$ 0.44 28.79 $\pm$ 4.70
Gemma-2-9b-it-FP8 71.89 $\pm$ 0.12 79.51 $\pm$ 1.48 32.56 $\pm$ 0.64 27.78 $\pm$ 6.06
Qwen2-72B-Instruct 82.17 $\pm$ 0.12 86.83 $\pm$ 1.10 37.40 $\pm$ 1.44 30.05 $\pm$ 4.55
Qwen2-72B-Instruct-FP8 81.99 $\pm$ 0.12 86.85 $\pm$ 0.31 37.66 $\pm$ 0.02 31.06 $\pm$ 2.04
Mistral-7B-Instruct-v0.3 59.12 $\pm$ 0.19 35.96 $\pm$ 6.41 9.01 $\pm$ 1.78 21.59 $\pm$ 0.25
Mistral-7B-Instruct-v0.3-FP8 58.70 $\pm$ 0.18 32.30 $\pm$ 0.00 7.62 $\pm$ 0.00 22.73 $\pm$ 0.00
@@
Attack 1: Quantization substitution.
We evaluate benchmark-based detection using four tasks: MMLU (Hendrycks et al., 2020), Math (Hendrycks et al., 2021), GPQA (Rein et al., 2024), and GSM8k (Cobbe et al., 2021). The MMLU benchmark is simulated 100 times by restricting the model’s output to multiple-choice tokens (e.g., “A”, “B”, etc.), computing their probabilities, and resampling proportionally after normalizing these probabilities to sum to 1. The GPQA and GSM8k tasks are evaluated using string matching with regular expressions, while the Math benchmark employs the math-verify library to verify answer correctness with these three tasks running 4 times. Our goal is to assess whether quantized models exhibit performance differences across these benchmarks and evaluation methods.
@@
Figure 2: Benchmark Accuracy (y-axis) vs. Temperature (x-axis) on different models.
As shown in Table 3, benchmark performance between quantized and original full-precision models is generally close, showing inherent randomness in evaluations and suggesting quantized models perform nearly indistinguishably from their full-precision counterparts. Auditors typically cannot repeat evaluations extensively due to rate limitations and budget constraints, making such evaluations more susceptible to quantized model substitution. Furthermore, benchmark evasion and randomized model substitutions can further shift results closer to the original model, allowing less capable quantized models to perform similarly. While usually sufficient, this further undermines benchmark verification effectiveness.
Attack 2: Hiding temperature parameter.
In Figure 2, we observe that for temperature $T\leq 1$, the standard deviation across most benchmarks is very small (especially for MMLU, where we have a larger sample size). Single-run benchmark results are fairly reliable; if a service provider claims to serve a state-of-the-art model, an auditor can easily detect substitution when observed performance falls slightly below published benchmark results. However, by not disclosing the temperature parameter used, the result could be attributed to either a higher temperature setting or an actual model substitution, making benchmark verification less effective.
4.1.5 Greedy decoding output
When given the option to do greedy decoding, users could theoretically compare API outputs directly with locally deployed models using identical parameters. However, our experiments with Llama-3-8B and Gemma-2-9B models showed that outputs never matched exactly between various API service and our local generation, across 100 UltraChat queries under greedy decoding. These discrepancies likely result from the non-deterministic of GPU computation and the algorithmic differences across different API providers. Therefore, exact output matching proves unreliable for detecting model substitution in practice. This results also align with the MMD difference from service provider shown in Gao et al. (2024).
4.2 Log probability verification
Method.
Log-probability comparison can be used for verifying if the underlying model is as claimed. When comparing the outputs of a deployed model against a known reference model, the log probabilities produced by the same model should be identical. Empirically, there exist small differences that can be attributed software and hardware difference. Thus, significant differences in log-probability outputs should indicate a model substitution.
Evaluation.
To evaluate the stability of log probabilities across different environments, we conducted testing using token-level log probabilities from greedy decoding on UltraChat Queries (Ding et al., 2023). We compared outputs across different versions of different inference frameworks (vLLM (Kwon et al., 2023) and Hugging Face Transformers), hardware configurations (H100 and A100 GPUs). In Figure 3, we show the log probabilities differences for the first 20 generated tokens of 2 sample generations. More samples can be found in Figure 4.
Weakness.
In Figure 3, we can observe that even in a controlled environment, logprobs still exhibit instability in early tokens across different software and hardware configurations. In practice, API providers serve multiple requests simultaneously, and different optimization techniques like continuous batching may also affect the results. As discussed in Section 4.1.5, the results from different API providers can even show token-level differences. In addition, this method cannot work on proprietary model since it requires white-box access to the reference model. Almost all API providers do not give the full logprobs access. If the full logprobs is provided, auditors may verify the served model by reconstructing the hidden dimension or output subspace, which is discussed in Appendix B.

### v2

@@ 4.1.1 Text classifier
Table 2: Binary classification accuracy between outputs of quantized and original models on UltraChat (Ding et al., 2023). Accuracies near 50% indicate failure to differentiate.
Attack: Quantization substitution. We test whether classifiers could distinguish full-precision from quantized variants of Llama-3.1-70B, Gemma-2-9B, Mistral-7B-v0.3, and Qwen2-72B. Using UltraChat prompts (temperature $0.6$), we train on 10k samples per variant (with 1k validation) and evaluate classifiers with BERT, T5, GPT-2, and LLM2Vec embeddings.
Analysis. Classification accuracies stay near chance across all model pairs and embedding backbones (Table 2), leaving us unable to reject $H_{0}$ that $M_{\text{spec}}$ and $M_{\text{alt}}$ produce indistinguishable outputs; deviations from 50% can be attributed to sampling noise. The result is consistent across families, parameter sizes, and INT8/FP8 variants, and persists under alternative train/validation splits. In contrast to Sun et al. (2025), who report around 98% cross-model and 80% cross-model size accuracy, the signal here is too weak to separate quantized from full-precision variants. We therefore find that text-based classifiers, even with strong encoders (BERT/T5/GPT-2/LLM2Vec), are ineffective for detecting quantization-based substitution under realistic sampling budgets.
4.1.2 Identity prompting
Method. Prompt the model to disclose its identity (e.g., “Who are you?”, “Who developed you?”) and string-match against known identifiers (family, developer), following the simple classifier setup in Huang et al. (2025).
Attack: Quantization substitution. We test whether quantization changes self-identification: for Llama3.1-70B-Instruct, Gemma2-9b-it, Mistral-7b-v3-Instruct, and Qwen2-72B-Instruct, we issue 500 identity queries per model/variant at temperature 0.6 and matched family (e.g., “Llama”) and developer (e.g., “Meta”).
Attack: System prompt override. In this setting, we explicitly set the system prompt to mislabel the model (e.g., “You are LLaMA, an AI assistant developed by Meta AI”) and then re-ran identity queries on Gemma2-9b-it, Mistral-7b-v3-Instruct, and Qwen2-72B-Instruct.
Analysis. The results for quantization substitution remain consistent: models typically disclose only broad identifiers such as the family or developer, and rarely reveal specific details like parameter count or quantization status. Queries asking for direct specifications are successful in less than 5% of cases, often resulting in refusals or incorrect responses from the model. When the system prompt is overridden, most models fail to correctly identify themselves, with accuracy dropping to nearly 0%. Instead, they often adopt the identity given by the system prompt, although some contradictions occasionally occur (e.g., restating the true developer). Overall, these findings indicate that identity prompting is ineffective for detecting substitutions within the same family and is highly sensitive to changes in the prompt, making it an unreliable method for detecting substitutions in practice.
4.1.3 Model equality testing
Method. When only text outputs are available, one can test whether two models induce the same distribution over completions by applying a Maximum Mean Discrepancy (MMD) test (Gao et al., 2024). Let $P$ and $Q$ denote the distributions of two candidate models,
$\mathrm{MMD}^{2}(P,Q)\;=\;\mathbb{E}_{x,x^{\prime}\sim P}[k(x,x^{\prime})]\;+\;\mathbb{E}_{x,x^{\prime}\sim Q}[k(x,x^{\prime})]\;-\;2\mathbb{E}_{x\sim P,x^{\prime}\sim Q}[k(x,x^{\prime})],$
with the Hamming kernel $k_{\mathrm{hamming}}(x,x^{\prime})=\sum_{i=1}^{L}\mathbf{1}\{x_{i}=x^{\prime}_{i}\}.$ We estimate significance by permutation testing: repeatedly shuffle completions between groups, recompute MMD, and reject $H_{0}$ if the observed statistic exceeds the $\alpha$-quantile of this null distribution.
Figure 1: Statistical power of MMD tests. (Left) Distinguishing quantized from original models when parameters are small. (Right) Power as a function of substitution probability $p$ for larger models, showing how mixtures of quantized and full-precision outputs affect detectability.
Attack: Quantization substitution. We first compare outputs from $M_{\text{spec}}$ (full precision) and $M_{\text{alt}}$ (INT8 quantized) models. Using Llama-3/3.1-70B, with $L=50$ and $T=1$, and 25 multilingual Wikipedia prompts, we conduct 100 Monte Carlo simulations with 1000 permutations per test. Power is computed as the probability of rejecting $H_{0}{:}P_{\text{actual}}{=}P_{\text{spec}}$ when the entire output distribution came from $M_{\text{alt}}$.
Attack: Randomized substitution. To simulate partial replacement, we construct mixtures of completions from $M_{\text{spec}}$ and $M_{\text{alt}}$ with substitution probability $p\in\{0.0,0.2,\dots,1.0\}$. We then repeat the same MMD test procedure to estimate how power changes as $p$ increases. This setup models a provider routing only a fraction of queries to the substituted model while serving the remainder from the specified one.
Analysis. Figure 1 shows that for smaller models (left panel), even full substitution by $M_{\text{alt}}$ remains statistically undetectable: power never rises beyond the significance level $\alpha$. For larger models (right panel), the test is more sensitive—power climbs with $p$ and reliably rejects $H_{0}$ when $p\geq 0.6$. However, for modest substitution ($p\leq 0.2$), rejection probability collapses to $\alpha$, making randomized substitution empirically indistinguishable from chance. These results imply that quantization substitution and partial randomized substitution are both viable evasion strategies, with the latter requiring only limited mixing to defeat auditing. Moreover, inference nondeterminism across providers can inflate MMD values even for identical models, risking false positives. As also noted by Gao et al. (2024), MMD-based auditing is only effective under strictly controlled local inference environments, limiting its practicality.
4.1.4 Benchmark-based detection
Method. An auditor may attempt benchmark-based detection (e.g., Chen et al. (2023)) by querying standardized benchmarks where baselines are well established. We evaluate four tasks: MMLU (Hendrycks et al., 2020), GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), and GPQA Diamond (Rein et al., 2024). At decoding temperature T=0.5T{=}0.5, we repeat evaluation 10 times and report mean accuracy and standard deviation (Table 3). MMLU is scored by restricting completions to multiple-choice tokens and averaging over 100 runs per query; free-form numeric tasks (GSM8K, MATH) are scored by regex matching or the math-verify library; GPQA is scored by multiple-choice accuracy.
Attack 1: Quantization substitution. We compare quantized and full-precision variants across all four benchmarks. As shown in Table 3, differences are small and typically within one standard deviation across repeated runs. This suggests that quantized models perform nearly indistinguishably from their full-precision counterparts, limiting the power of benchmark-only detection.
@@ 4.1.4 Benchmark-based detection
Meta-Llama-3-8B-Instruct 62.69 $\pm$ 0.18 61.14 $\pm$ 3.47 20.65 $\pm$ 3.43 22.62 $\pm$ 0.22
Meta-Llama-3-8B-Instruct-FP8 62.43 $\pm$ 0.26 60.90 $\pm$ 4.10 14.91 $\pm$ 2.49 20.14 $\pm$ 0.24
Meta-Llama-3-70B-Instruct 78.05 $\pm$ 0.08 88.06 $\pm$ 1.44 35.69 $\pm$ 1.33 29.60 $\pm$ 0.51
Meta-Llama-3-70B-Instruct-FP8 77.88 $\pm$ 0.13 87.35 $\pm$ 1.24 35.75 $\pm$ 1.16 33.12 $\pm$ 0.30
Gemma-2-9b-it 71.86 $\pm$ 0.08 81.80 $\pm$ 1.35 33.41 $\pm$ 0.28 28.64 $\pm$ 2.97
Gemma-2-9b-it-FP8 71.92 $\pm$ 0.11 79.41 $\pm$ 1.14 32.53 $\pm$ 0.34 27.81 $\pm$ 3.22
Qwen2-72B-Instruct 82.18 $\pm$ 0.08 86.72 $\pm$ 1.00 37.39 $\pm$ 1.41 29.93 $\pm$ 2.71
Qwen2-72B-Instruct-FP8 81.98 $\pm$ 0.08 86.82 $\pm$ 0.97 37.67 $\pm$ 1.39 31.08 $\pm$ 1.96
Mistral-7B-Instruct-v0.3 59.15 $\pm$ 0.10 35.90 $\pm$ 4.54 8.94 $\pm$ 1.24 21.60 $\pm$ 0.17
Mistral-7B-Instruct-v0.3-FP8 58.77 $\pm$ 0.13 32.20 $\pm$ 4.02 7.68 $\pm$ 1.23 22.72 $\pm$ 0.19
@@
Attack 2: Hiding temperature parameter. We also evaluate benchmark accuracy under varying decoding temperature (Figure 2). For $T\leq 1$, accuracy variance across runs is low, and single-run results are stable. However, if the provider does not disclose or allow control of temperature, observed accuracy drops could be misattributed—either to hidden parameter choices or to substitution—reducing audit reliability.
@@
Figure 2: Benchmark accuracy versus decoding temperature across models and tasks.
Analysis. Figure 2 shows that benchmark accuracy degrades with higher $T$, and the size of this drop is comparable to quantization-induced differences. This means auditors testing $H_{0}{:}P_{\text{actual}}{=}P_{\text{spec}}$ cannot distinguish whether deviations from baseline arise from $M_{\text{alt}}$ substitution or from hidden parameter settings. Caching further complicates detection: if providers cache outputs for known benchmark queries, apparent accuracy may remain aligned with $M_{\text{spec}}$ even under substitution, suppressing deviations. Overall, benchmark-based detection can expose large-scale substitution, but smaller changes (e.g., quantization) are hidden by natural variability, and confounds like temperature or caching make $H_{0}$ rejection unreliable in practice.
4.1.5 Greedy decoding outputs
Method. We test whether greedy decoding, which removes sampling randomness, yields reproducible completions across settings. Beyond substitution attacks, we also investigate how greedy decoding behaves in real-world auditing conditions, where inference is accessed only through APIs. This lets us assess not just sensitivity to quantization or fine-tuning, but also whether greedy decoding (and text-based verification more broadly) is robust when applied in deployment settings. Using 50 UltraChat queries, we compare local inference runs (Transformers 4.40.0 on H100 GPUs) against API providers’ outputs on the same Gemma-2-9B model. We also assess self-consistency across repeated API calls. Agreement is measured by exact match of the first $k$ tokens, with $k=20$ and $k=100$.
Attack: Quantization and fine-tuned substitution. For Llama-3-8B, we compare greedy outputs from the reference model against its quantized (FP8) and fine-tuned variants (e.g., models trained for other context length or on domain-specific corpora). While local inference across frameworks is largely consistent, both quantization and fine-tuning introduce significant token-level discrepancies, reducing agreement with the baseline sequence. This shows that greedy decoding can, in theory, detect within-family substitutions when the environment is controlled.
Figure 3: Greedy decoding agreement rate with local baseline (Transformers 4.40.0, H100). Left: Gemma-2-9B APIs across providers. Right: Llama-3-8B variants including quantized and fine-tuned models.
Analysis. Figure 3 shows that greedy decoding does not ensure reproducibility: even with sampling disabled, agreement across provider APIs for Gemma-2-9B is below 5%, and agreement for fine-tuned or quantized Llama-3-8B variants drops well below the baseline. Under $H_{0}:P_{\text{actual}}{=}P_{\text{spec}}$, greedy outputs should match exactly, but observed divergence is often larger than that induced by quantization or fine-tuning. This reflects inference nondeterminism (He & Lab, 2025), whose sources (e.g., backend variability in kernels, tokenization, or caching) are discussed in the next section. The key implication is that while greedy decoding can detect substitution in a controlled local environment, it fails to do so in realistic auditing settings where backend nondeterminism dominates, making $M_{\text{spec}}$ and $M_{\text{alt}}$ indistinguishable.
4.1.6 Log probability verification
Method. In theory, logprobs comparison can be used to verify whether the served model matches a claimed reference: per-token log probabilities should align closely. In practice, the reliability of this approach depends on the magnitude of nondeterminism affecting the logprobs.
Evaluation. To assess stability, we compare token-level log probabilities from greedy decoding on UltraChat queries, across multiple inference frameworks (vLLM (Kwon et al., 2023) vs. Hugging Face Transformers (Wolf et al., 2020)), GPU types (H100, A100), and software versions. Figure 4 shows that even for the same model, logprob traces for the first 20 tokens can diverge across stacks. On the subset of completions where the generated tokens agreed, we further compare the magnitude of the assigned log probabilities and find minor variations across environments. Additional examples are given in Appendix B.
Figure 4: Log probabilities for the first 20 shared tokens under greedy decoding for UltraChat queries across frameworks/hardware.
Analysis and weakness. Logprob traces often appear nearly identical across token positions in controlled settings, and the magnitude differences are still distinguishable when compared to quantized or fine-tuned variants. In production, however, batching and heterogeneous backends introduce much larger variance. Recent work (He & Lab, 2025) shows that a key source of this instability is the lack of batch invariance in kernels such as RMSNorm, matrix multiplication, and attention: the same request may yield different results depending on how many other requests are batched concurrently or how prefill/chunking is scheduled. Even at $T=0$, greedy decoding can diverge after only tens of tokens, making logprob distributions very noisy. Prototype batch-invariant kernels exist but incur throughput costs, so current APIs typically remain nondeterministic. For auditors, this means logprob-based detection is vulnerable to false alarms unless requests are isolated or providers explicitly expose batch-invariant execution. In addition, this method requires a reference instance of $M_{\text{spec}}$ and API access to log probabilities. Stronger auditing methods that attempt to reconstruct embedding subspaces (Appendix B) are generally infeasible under current API non-disclosure of full logprobs.
4.2 Internal Activations Verification
Method. TopLoc (Ong et al., 2025) introduces a locality-sensitive hashing scheme over intermediate activations to produce compact proofs of correct execution. Here, locality-sensitive hashing (LSH) refers to mapping high-dimensional activation tensors into small fingerprints such that similar activations yield similar hashes, while even small unauthorized changes cause detectable differences. Instead of storing full tensors, TopLoc extracts the top-$k$ largest activation values and indices and encodes them as a polynomial congruence, yielding proofs that are robust to GPU nondeterminism and algebraic reorderings. This allows compact verification (about 100 bytes per tens of tokens) while maintaining accurate detection under adversarial substitution.
Analysis and weakness. TopLoc can detect most model substitutions (e.g., quantized variants, fine-tuned models) even under different software and hardware settings, aligning with the behaviors observed in logprobs testing. However, verification still requires the auditor to recompute the same hidden activations, i.e., to run $M_{\text{spec}}$ locally or rely on an attested recomputation (e.g., a TEE), so it is infeasible without attestation for proprietary models. In addition, batching variance and speculative decoding could still remain problematic, as both can introduce enough instability in the forward pass or generated tokens to disrupt verification.